OpenStudy (pflori1234_pop):

Find the ratio of the gravitational force between two planets if the masses of both planets are tripled but the distance between them stays the same. Express your answer as an integer. F^2/F^1

OpenStudy (pflori1234_pop):

@TheSmartOne

TheSmartOne (thesmartone):

\(\Large \sf f_{gav} = G \cdot \frac{m_1 \times m_2}{distance}\) m1 and m2 are the two masses G is a constant, distance will remain the same hint: substitute \(3m_1\) in place of \(m_1\) and \(3m_2\) in place of \(m_2\)

OpenStudy (pflori1234_pop):

so 9? @TheSmartOne

OpenStudy (pflori1234_pop):

Thank you so much!

TheSmartOne (thesmartone):

That's correct, I deleted my previous reply because I accidentally got scared that it was wrong and that there was a square, but upon reviewing the formula, I got myself all together Great job! :D